# Supplementary material – Approximate Dynamic Programming for Constrained Piecewise Affine Systems with Stability and Safety Guarantees

### v2

@@ VI-B Example 2: centralized adaptive cruise control
Methods
CPU time (mean/max)
Total cost
Safety rate
MPC with $N=4$
0.1592 / 0.4593
21.7664
100%
MPC with $N=5$
0.2850 / 0.7703
21.6653
100%
MPC with $N=6$
0.4630 / 1.4619
21.5997
100%
$\hat{\pi}^{\mathrm{im}}$ with $M=15$
0.0605 / 0.1036
23.0097
98%
$\hat{\pi}^{\mathrm{im}}$ with $M=20$
0.0916 / 0.1621
22.6613
100%
$\hat{\pi}^{\mathrm{im}}$ with $M=25$
0.1372 / 0.2991
22.3042
100%
$\hat{\pi}^{\mathrm{im}}$ with $M=30$
0.5487 / 1.3904
22.0716
100%
$\hat{\pi}^{\mathrm{ex}}_{\mathrm{proj}}$ with $M=15$
0.0008 / 0.0028
23.4043
99%
$\hat{\pi}^{\mathrm{ex}}_{\mathrm{proj}}$ with $M=20$
0.0007 / 0.0027
22.9568
99%
$\hat{\pi}^{\mathrm{ex}}_{\mathrm{proj}}$ with $M=25$
0.0007 / 0.0027
22.7022
99%
$\hat{\pi}^{\mathrm{ex}}_{\mathrm{proj}}$ with $M=30$
0.0008 / 0.0030
22.7562
99%
RL in [15] with $M=30$,$N=5$
0.0809 / 0.3570
29.7561
85%
@@ A-A3 Equivalence of the VI sequences for options 1 and 2
In option 2, by iterating $J_{k}^{\mathrm{soft}}(x)=\Gamma_{\mathrm{p,2}}J_{k-1}^{\mathrm{soft}}(x),\;\forall x\in\mathcal{X}$ from $k$ to 0, we can get the expression of $J_{k}^{\mathrm{soft}}(\cdot)$ via the batch approach:
@@
From (29), we notice that the difference between the value functions in options 1 and 2 at the same iteration is $P(x,X)$, which is always continuous in $x$ and equals zero if $x\in X$. Combining (29) with the first and second parts of the proof proves the statements of Theorem 1.
@@ A-C PWA functions and their mixed-integer formulations
Mixed-logical dynamical formulation of PWA systems: In the traditional MPC framework for PWA systems [5], the PWA model is usually converted into an equivalent mixed-logical dynamical form, so that the MPC problem is formulated as a solvable MILP problem. By specifying the partition $\left\{\mathcal{C}_{i}\right\}_{i=1}^{s}$ as $\{{\mathcal{C}}_{i}\}_{i=1}^{s}\triangleq\{\{\left[\begin{array}[]{ll}x^{T}&u^{T}\end{array}\right]^{T}\in\mathbb{R}^{n_{x}+n_{u}}:{\Psi^{(x)}_{i}}x+{\Psi^{(u)}_{i}}u\leq{\psi_{i}}\}\}^{s}_{i=1}$, according to [1], (1) can always be described as a mixed-logical dynamical (MLD) system. An equivalent representation [57] is given by
@@
where $M$, $m$, and $M_{i}^{*}$ are some bounds of affine functions of $x$ and $u$ over $X\times U$, and $\Delta_{i}\in\{0,1\},i=1,2,\dots,s$ are binary variables. The detailed definitions of the bounds $M$, $m$, and $M_{i}^{*}$ can be found in [57]. The last inequality in (A-C) is used to formulate the logic $[x^{T}\;\;u^{T}]^{T}\in\cup^{s}_{i=1}\mathcal{C}_{i}$. This can be generalized to convert any UoP constraint into a mixed-integer constraint, such as $x_{0}\in X_{\mathrm{in}}$ in (V-C). Mixed-integer formulation of penalty functions in (III): As (12) and (14) are all minimization problems, the maximum operation in $P(\cdot,\cdot)$ can be eliminated by introducing some linear inequalities and continuous variables. On the other hand, the minumum operation in $P(\cdot,\cdot)$ results in a mixed-integer formulation similar to (A-C).

### v3

@@ VI-B Example 2: centralized adaptive cruise control
Methods CPU time (mean/max) Total cost Safety rate
MPC with $N=4$ 0.1592 / 0.4593 21.7664 100%
MPC with $N=5$ 0.2850 / 0.7703 21.6653 100%
MPC with $N=6$ 0.4630 / 1.4619 21.5997 100%
$\hat{\pi}^{\mathrm{im}}$ with $M=15$ 0.0605 / 0.1036 23.0097 98%
$\hat{\pi}^{\mathrm{im}}$ with $M=20$ 0.0916 / 0.1621 22.6613 100%
$\hat{\pi}^{\mathrm{im}}$ with $M=25$ 0.1372 / 0.2991 22.3042 100%
$\hat{\pi}^{\mathrm{im}}$ with $M=30$ 0.5487 / 1.3904 22.0716 100%
$\hat{\pi}^{\mathrm{ex}}_{\mathrm{proj}}$ with $M=15$ 0.0008 / 0.0028 23.4043 99%
$\hat{\pi}^{\mathrm{ex}}_{\mathrm{proj}}$ with $M=20$ 0.0007 / 0.0027 22.9568 99%
$\hat{\pi}^{\mathrm{ex}}_{\mathrm{proj}}$ with $M=25$ 0.0007 / 0.0027 22.7022 99%
$\hat{\pi}^{\mathrm{ex}}_{\mathrm{proj}}$ with $M=30$ 0.0008 / 0.0030 22.7562 99%
RL in [15] with $M=30$,$N=5$ 0.0809 / 0.3570 29.7561 85%
@@ A-A3 Equivalence of the VI sequences for options 1 and 2
In option 2, by iterating $J_{i}^{\mathrm{soft}}(x)=\Gamma_{\mathrm{p,2}}J_{i-1}^{\mathrm{soft}}(x),\;\forall x\in\mathcal{X}$ from $i=1$ to $i=k$, we can get the expression of $J_{k}^{\mathrm{soft}}(\cdot)$ via the batch approach:
${J_{k}^{\mathrm{soft}}}\left({{x}}\right)$
$=$ $\min_{u_{0}\in U}\;l(x,u_{0})+J_{k-1}^{\mathrm{soft}}\left(f_{\mathrm{PWA}}(x,u_{0})\right)+P_{k-1}\left(f_{\mathrm{PWA}}(x,u_{0})\right)$
$=$ $\min_{u_{0},u_{1},x_{0},x_{1}}l(x_{0},u_{0})+l(x_{1},u_{1})+J_{k-2}^{\mathrm{soft}}\left(f_{\mathrm{PWA}}(x_{1},u_{1})\right)$
$\quad\quad\quad\quad+P_{k-1}\left(x_{1}\right)+P_{k-2}\left(f_{\mathrm{PWA}}(x_{1},u_{1})\right)$
$\quad\mathrm{s.t.}\;x_{0}=x,\;x_{1}=f_{\mathrm{PWA}}(x_{0},u_{0}),u_{0},u_{1}\in U$
@@
$=$ $\min_{\{u_{i}\}^{k}_{i=0},\;\{x_{i}\}^{k}_{i=0}}\sum_{i=0}^{k-1}\left(l\left(x_{i},u_{i}\right)+P_{k-1-i}\left(x_{i+1}\right)\right)+J_{0}^{\mathrm{soft}}(x_{k})$
[…]
$\quad\quad u_{i}\in U,i=0,\ldots,k-1.$ (29)
Since $P_{0}(x)=0,\forall x\in\mathcal{X}$ and $P_{k}(x)=P(x,X),\forall x\in\mathcal{X}$ and $\forall k>0$, by noticing that $l_{\mathrm{p}}(x,u)=l(x,u)+P(x,X)$ we have
${J_{k}^{\mathrm{soft}}}\left({{x}}\right)=$ $\min_{\{u_{i}\}^{k}_{i=0},\;\{x_{i}\}^{k}_{i=0}}\sum_{i=0}^{k-1}l_{\mathrm{p}}\left(x_{i},u_{i}\right)+J_{0}^{\mathrm{soft}}(x_{k})-P(x,X)$
[…]
$\quad\quad u_{i}\in U,i=0,\ldots,k-1.$ (30)
By comparing the right-hand side of (A-A3) and (28), we can conclude that
${J_{k}^{\mathrm{soft}}}\left({{x}}\right)=\mathrm{the\;optimal\;value\;of}\;(28)-P(x,X),\;\mathrm{for}\;k=0,1,...$ (31)
From (31), we notice that the difference between the value functions in options 1 and 2 at the same iteration is $P(x,X)$, which is always continuous in $x$ and equals zero if $x\in X$. Combining (31) with the first and second parts of the proof proves the statements of the theorem.
@@ A-C PWA functions and their mixed-integer formulations
Mixed-logical dynamical formulation of PWA systems: In the traditional MPC framework for PWA systems [5], the PWA model is usually converted into an equivalent mixed-logical dynamical form, so that the MPC problem is formulated as a solvable MILP problem.
By specifying the partition $\left\{\mathcal{C}_{i}\right\}_{i=1}^{s}$ as $\{{\mathcal{C}}_{i}\}_{i=1}^{s}\triangleq\{\{\left[\begin{array}[]{ll}x^{T}&u^{T}\end{array}\right]^{T}\in\mathbb{R}^{n_{x}+n_{u}}:{\Psi^{(x)}_{i}}x+{\Psi^{(u)}_{i}}u\leq{\psi_{i}}\}\}^{s}_{i=1}$, according to [1], (1) can always be described as a mixed-logical dynamical (MLD) system. An equivalent representation [57] is given by
@@
where $M$, $m$, and $M_{i}^{*}$ are some bounds of affine functions of $x$ and $u$ over $X\times U$, and $\Delta_{i}\in\{0,1\},i=1,2,\dots,s$ are binary variables. The detailed definitions of the bounds $M$, $m$, and $M_{i}^{*}$ can be found in [57].
The last inequality in (A-C) is used to formulate the logic $[x^{T}\;\;u^{T}]^{T}\in\cup^{s}_{i=1}\mathcal{C}_{i}$. This can be generalized to convert any UoP constraint into a mixed-integer constraint, such as $x_{0}\in X_{\mathrm{in}}$ in (V-C).
Mixed-integer formulation of penalty functions in (III): As (12) and (14) are all minimization problems, the maximum operation in $P(\cdot,\cdot)$ can be eliminated by introducing some linear inequalities and continuous variables. On the other hand, the minumum operation in $P(\cdot,\cdot)$ results in a mixed-integer formulation similar to (A-C).